# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ 5.3 Generalized impulse response functions
The effects of the shocks in our structural LSTVAR model may depend on the initial state of the economy as well as on the sign and size of the shock, which makes the conventional way of calculating impulse responses unsuitable. Therefore, we consider the general impulse response function (GIRF) (Koop et al., 1996) that accommodates such features and is defined as
@@
where $h$ is the horizon. The first term on the right side of (5.1) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{i}\in\mathbb{R}$ in the $i$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the $i$th structural shock of sign and size $\delta_{i}$ arrives at time $t$ as opposed to all shocks being random. The GIRFs also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
@@
Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$. To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives. Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant (there are $366$ such histories for the Low EPU Regime and $80$ for the High EPU Regime). For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B. Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
@@
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of the High EPU Regime are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating. All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points. Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
@@
The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf. Lanne and Virolainen, 2025). The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in the Low EPU Regime and decreases in the High EPU Regime. A possible explanation for the decrease in the EPUI in the High EPU Regime could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes. Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al. (2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital. Although many of the GIRFs exhibit a short-term increase of IPI in the High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
@@
In both regimes, a positive CPU shock increases consumer prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023). Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup. However, we find that the inflationary effects are stronger in the High EPU Regime than in the Low EPU Regime. One possible explanation is that the firms are more risk averse in the periods of high EPU, and thereby react more sensitively. Producer prices, in turn, mostly decrease in the Low EPU Regime but increase in the High EPU Regime.
@@
The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases in both regimes, but clearly more so in the High EPU Regime. Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of the High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU. Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.
@@ 6 Conclusion
Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017). We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well. Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
@@
In addition to establishing the identification of the shocks, we have discussed the problem of labelling them, and employed blended identification (cf. Carriero et al., 2024) to address weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices. We have proposed estimating the model parameters by the method of penalized maximum likelihood, and introduced a three-step estimation procedure by adapting the approach of Lanne et al. (2017) to nonlinear SVAR models. Building on the results of Saikkonen (2008), we have also provided a sufficient condition for ergodic stationarity of our structural STVAR model. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@
In an empirical application to U.S. data from 1987:4 to 2024:12, we have studied the macroeconomic effects of the climate policy uncertainty shock using a two-regime logistic STVAR model. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. To distinguish between periods low and high economic policy uncertainty, we used the first lag of the EPU index (Baker et al., 2016) as the switching variable. Our results are in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024). However, we found that while a positive CPU shock decreases production and increases inflation in times of both low and high EPU, its inflationary effects are stronger during the periods of high EPU.
Appendix A Stationarity and ergodicity
@@
Establishing stationarity and ergodicity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator. Saikkonen (2008) derives a sufficient condition for ergodic stationarity of the smooth transition vector error correction (STVEC) model. Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008). However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes. As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model is not nested to the model of Saikkonen (2008). The applicability of the results of Saikkonen (2008) also requires that the weights are not exogenous but functions of the preceding $p$ observations (the time $t$ weights are allowed to depend on the time $t$ observation of an IID process independent of $e_{t}$, but we do not discuss this option further, as it violates our assumption of $\mathcal{F}_{t-1}$-measurable or nonrandom weights). Nonetheless, we will show in this section that the stationarity condition applies to our model as well.
@@
$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (A.1)
@@
$\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (A.2)
@@
    The transition weight $\alpha_{m,t}$ is an $\mathcal{F}_{t-1}$-measurable function of $\{y_{t-j},j=1,...,p\}$ for all $t$ and $m=1,...,M$.
Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (10) of Saikkonen (2008). Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the process (to the best of our knowledge, there does not currently exist results presenting a sufficient necessary condition for STVAR models). Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example. Finally, Condition (III) ensures that the assumptions for the transition function imposed by Saikkonen (2008) are satisfied, ruling out nonrandom exogenous weights (which are accommodated by our identification results).
The following theorem, which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
Theorem A.1.
Consider the STVAR process $y_{t}$ defined in Equations (3.2) and (3.3), and suppose Assumptions 1 and A.1 are satisfied. Then, almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, the process $\boldsymbol{y}_{t}=(y_{t},...,y_{t-p+1})$ is a $(1+||x||^{2})$-geometrically ergodic Markov chain. Thus, there exists a choice of initial values $y_{-p+1},...,y_{0}$ such that the process $y_{t}$ is strictly stationary, second-order stationary, and $\beta$-mixing with geometrically decaying mixing numbers.
Theorem A.1 is proven in Appendix A.1. A notable difference of Theorem A.1 to Theorem 1 of Saikkonen (2008) (and Theorem 1 of Kheifets and Saikkonen, 2020) is that our results holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, while the result of Saikkonen (2008) holds everywhere. To be more specific, the result of Theorem A.1 holds with such values of $B_{1},...,B_{M}$ that the resulting impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ is invertible for all $t$, and Lemma 1 shows that this holds everywhere expect in regions of $\mathbb{R}^{d\times dM}$ that have a Lebesgue measure zero.
Assumption A.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption A.1(I) is computationally demanding to verify in practice (e.g., Chang and Blondel, 2013). This makes it poorly suitable for restricting the parameter space in the numerical estimation of parameters discussed in Section 4. During estimation, this condition would have to be verified repeatedly a very large of times, making the estimation extremely tedious. Hence, it is useful make use of the following condition in the estimation, which is necessary for Assumption A.1(I).
Condition A.1.
@@
where $\rho(\boldsymbol{A}_{m})$ is the spectral radius of $\boldsymbol{A}_{m}$, $m=1,...,M$.
Condition A.1 states that the usual stability condition is satisfied by each of the regimes. It is necessary for Assumption A.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). After estimation, Assumption A.1(I) can be checked for the solutions of interest.
Several methods have been proposed for bounding the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
Appendix B Proofs
B.1 Proof of Lemma 1

### v5

@@ 5.3 Impulse response analysis
The effects of the shocks in our structural LSTVAR model may depend on the initial state of the economy as well as on the sign and size of the shock, making the conventional way of calculating impulse responses unsuitable. Therefore, we consider the generalized impulse response function (GIRF) (Koop et al., 1996) that accommodates such features, defined as
@@
where $h$ is the horizon. The first term on the right side of (5.1) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{i}\in\mathbb{R}$ in the $i$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the $i$th structural shock of sign and size $\delta_{i}$ arrives at time $t$ as opposed to all shocks being random. Since our model has the $p$-step Markov property, the conditioning set $\mathcal{F}_{t-1}$ can be replaced by the lag vector $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$. The GIRF also facilitates tracing out the effects of the shocks on the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
@@
To study the state-dependent effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives. Specifically, for each Regime $m\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{m,t}$ is greater than $0.75$, indicating that Regime $m$ is clearly dominant. For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B. Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
@@
Figure 2: Generalized impulse response functions to the identified CPU shock h=0,1,.,36h=0,1,....,36 months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, RATE, and the transition weights of the High EPU Regime are depicted in each row. The responses of IPI and CPI are accumulated to ($100\times$)log-levels. The left column shows the responses conditionally on the Low EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on the High EPU Regime dominating. All the GIRFs have been scaled so that the instantaneous increase in CPUI is 55 points. Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
@@
Figure 2 illustrates the distribution of the GIRFs in each regime in a so-called ”shotgun plot”, which depicts each GIRF using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf. Lanne and Virolainen, 2025). The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs. EPUI increases in both regimes, but response is stronger in the High EPU Regime.
Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes. Huang and Punzi (2024) attribute the decline in output to firms postponing investment, whereas Fried et al. (2022) and Khalil and Strobel (2023) emphasize reallocation of capital away from carbon-intensive sectors. We find that the contraction is slightly stronger in the High EPU Regime than in the Low EPU Regime, possibly reflecting greater risk aversion among firms and financial intermediaries during periods of elevated economic policy uncertainty.
@@
In both regimes, a positive CPU shock increases prices, in line with Huang and Punzi (2024) and Khalil and Strobel (2023). Huang and Punzi (2024) find that a positive CPU shock increases inflation by acting as a negative supply shock, whereas according to Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices preemptively to lower the probability of being stuck with a negative markup. However, we find that the inflationary effects are stronger in the High EPU Regime than in the Low EPU Regime. One possible explanation is that the firms might be more risk averse in the periods of high EPU and thereby react more sensitively.
@@
The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases in both regimes, but clearly more so in the High EPU Regime. Finally, consistent with the GIRFs of EPUI, a positive CPU shock increases the transition weights of the High EPU Regime, more so in the High EPU Regime, where there is first a short-term decline in the weights. In other words, a positive CPU shock generally drives the economy towards the High EPU Regime. Overall, our results suggest that a positive CPU shock reduces production and raises inflation in both regimes, but its effects, particularly on inflation, are stronger under high economic policy uncertainty.
@@ 6 Conclusion
Linear structural vector autoregressive models can be identified statistically by non-Gaussianity, provided that the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017). This paper is the first to extend identification by non-Gaussianity to nonlinear SVAR models. Specifically, we have shown that structural smooth transition vector autoregressive models, featuring a time-varying impact matrix defined as a weighted sum of regime-specific impact matrices, are likewise identified under these assumptions. While we have focused on exogenous, logistic, and threshold transition weights typically used in macroeconomic applications, our results can be extended to other suitable weight functions as well.
@@
In addition to establishing the identification of the shocks, we have discussed the problem of labeling them. We have also employed blended identification (cf. Carriero et al., 2024) to address weak identification, which in our experience is often but not exclusively related to the ordering and signs of the columns of the regime-specific impact matrices. We have proposed estimating the model parameters by the method of penalized maximum likelihood, and we introduced a three-step estimation procedure by adapting the approach of Lanne et al. (2017) to nonlinear SVAR models. Building on the results of Saikkonen (2008), we have also provided a sufficient condition for ergodic stationarity of our structural STVAR model. The introduced methods are implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available via the CRAN repository.
@@
In an empirical application to U.S. data from 1987:4 to 2024:12, we have studied the macroeconomic effects of the climate policy uncertainty shock using a two-regime logistic STVAR model. As a measure of climate policy uncertainty, we employ the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. To distinguish between periods of low and high economic policy uncertainty, we use the first lag of the EPU index (Baker et al., 2016) as the switching variable. Our results are in line with the previous literature suggesting that a positive CPU shock reduces production and raises inflation (Khalil and Strobel, 2023 and Huang and Punzi, 2024), while Fried et al. (2022) found that it decreases production. However, we found that while a positive CPU shock reduces production and raises inflation in times of both low and high EPU, its effects, particularly on inflation, are stronger during the periods of high EPU.
Appendix A Identification of the shocks in structural TVAR models
The threshold VAR (TVAR) model of Tsay (1998) is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
$\alpha_{m,t}=\left\{\begin{matrix}1&\text{if}\ \ r_{m-1}<z_{t}\leq r_{m},\\ 0&\text{otherwise},\phantom{aaaaaaa}\end{matrix}\right$. (A.1)
where $r_{m}\in\mathbb{R}$, m=1,,M1m=1,...,M-1, are the threshold parameters that satisfy $\infty<r_{1}<...<r_{M-1}<\infty$, $r_{0}\equiv$-$\infty$, $r_{M}\equiv\infty$, and $z_{t}\in\{y_{it-j},i=1,...,d,j=1,...,p\}$ is the switching variable, which we assume the be a lagged endogenous variable (up to the lag $p$). At each tt‚ the model defined in Equations (2.1), (2.2) and (A.1) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$. Notably, the TVAR model can be obtained as special case of the LSTVAR model when the smoothing parameter tends to infinity, i.e., $\gamma=\infty$ in (2.3). Also, exogenous transition weights may sometimes be binary or close-to-binary. Therefore, it is useful to establish the identification result also in such special cases of the STVAR model.
Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and m=1,,Mm=1,...,M, specifically, either exogenous or of the threshold form (A.1). Then, at each $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$. If the allocation of the time periods to the regimes is uniquely identified, it then follows from Lemma 2 that $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns. Thus, to obtain identification up to ordering and signs, identification of the threshold and AR parameters is established in the following proposition, which is proven in Appendix C.4.
Proposition A.1.
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 satisfied; and $\alpha_{m,t}$ either of the threshold form (A.1) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and m=1,,Mm=1,...,M. Suppose the following conditions hold:
  1. (A)
    $(\phi_{m},\text{vec}(A_{m,1}),...,\text{vec}(A_{m,p}))\neq(\phi_{n},\text{vec}(A_{n,1}),...,\text{vec}(A_{n,p}))$ for all $m\neq n\in\{1,...,M\}$,
  2. (B)
    (for exogenous weights) there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, i=1,,Mi=1,...,M, are linearly independent.
Then, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, m=1,,Mm=1,...,M, and (for threshold models) $r_{1},...,r_{M-1}$ are uniquely identified. Moreover, $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns.
Condition (A) of Proposition A.1 implies distinguishable regimes, whereas Condition (B) guarantees (for exogenous weights) that each regime prevails in at least one time period. Under these conditions, the AR parameters and the partition of the time periods into the regimes are uniquely identified. Moreover, the regime-specific impact matrices $B_{1},...,B_{M}$ are identified up to ordering and signs of their columns. [^10] The identification is for every invertible $B_{1},...,B_{M}$ and not just for almost every, as $B_{y,t}=B_{m}$ for some $m$ for all $t$, implying that $B_{y,t}$ is invertible for all invertible $B_{m}$. Thus, the result of Lemma 2 applies for all invertible $B_{1},...,B_{M}$.
By Proposition A.1, the shocks are identified by fixing the ordering and signs of the columns of $B_{1},...,B_{M}$. In general, fixing them in one regime does not necessarily determine them in the others, and whether such constraints are purely normalizations or overidentifying depends on the distributions of the shocks. In practice, we recommend fixing the ordering and signs in all $B_{1},...,B_{M}$ as a part of blended identification in Section 3.3, which combines non-Gaussianity with supplementary information to address weak identification discussed in Sections 3.2 and 3.3.
Appendix B Stationarity and ergodicity
@@
Establishing ergodic stationarity of the model is a common practice in time series econometrics, as it facilitates obtaining desirable asymptotic properties such as consistency of the estimator. Saikkonen (2008) derives a sufficient condition for ergodic stationarity of the smooth transition vector error correction (STVEC) model. Kheifets and Saikkonen (2020) make use of the results of Saikkonen (2008) to show that the stationarity condition readily applies for a STVAR model that can be obtained as a special case of the STVEC model of Saikkonen (2008). However, in both Saikkonen (2008) and Kheifets and Saikkonen (2020), the conditional covariance matrix of the process is defined as a weighted sum of positive definite error term covariance matrices of the regimes. As discussed in Section 3.1, the conditional covariance matrix (3.4) of our structural STVAR model has a different form, and therefore, our model is not nested to the model of Saikkonen (2008). Nonetheless, we show in this section that the stationarity condition of Saikkonen (2008) applies to our model as well, given that the transition weights are either logistic or of the threshold form, with an extra assumption required for the logistic weights.
@@
$\rho(\mathcal{A})=\underset{j\rightarrow\infty}{\limsup}\left(\underset{A\in\mathcal{A}^{j}}{\sup}\rho(A)\right)^{1/j},$ (B.1)
@@
$\boldsymbol{A}_{m}=\underset{(dp\times dp)}{\begin{bmatrix}A_{m,1}&A_{m,2}&\cdots&A_{m,p-1}&A_{m,p}\\ I_{d}&0&\cdots&0&0\\ 0&I_{d}&&0&0\\ \vdots&&\ddots&\vdots&\vdots\\ 0&0&\ldots&I_{d}&0\end{bmatrix}},\ m=1,...,M.$ (B.2)
@@
    The transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, follow either the threshold (A.1) or logistic (2.3) process with an endogenous switching variable $z_{t}\in\{y_{it-j},i=1,...,d,j=1,...,p\}$, and
  4. (IV)
    for logistic weights (with $M=2$ assumed), the matrix $B_{1}^{-1}B_{2}$ has no negative real eigenvalues.
Condition (I) states that the JSR of the companion form AR matrices of the regimes is strictly less than one, and it is analogous to Condition (19) of Saikkonen (2008). Note that this condition is sufficient but not necessarily necessary for ergodic stationarity of the model. Condition (II) is analogous Assumption 1 of Saikkonen (2008), and it is innocuous in practice but rules out bounded error distributions, for example. Condition (III), in turn, restricts the form of the regime weights to either the standard threshold or logistic processes with endogenous switching, ensuring piecewise constancy or continuity of the weights. Finally, Condition (IV) rules out singular convex combinations of the regime-specific impact matrices in the two-regime logistic case, which guarantees positive definiteness of the conditional covariance matrix for all values of the transition weights. In the threshold case, positive definiteness is automatically guaranteed since the impact matrix coincides with one of the regime-specific matrices, each assumed invertible.
The following theorem (proven in Appendix C.5), which is analogous to Theorem 1 in Saikkonen (2008) and Theorem 1 in Kheifets and Saikkonen (2020), states the results.
Theorem B.1.
[…]
Assumption B.1 is sufficient for ergodic stationarity of our structural STVAR model, but Assumption B.1(I) is computationally demanding to verify in practice (see, e.g., Chang and Blondel, 2013). This makes it poorly suitable for restricting the parameter space in the numerical estimation of parameters discussed in Section 4. During estimation, this condition would have to be verified repeatedly a very large number of times, making the estimation extremely tedious. Hence, it is useful to make use of the following condition in the estimation, which is necessary for Assumption B.1(I).
Condition B.1.
@@
where $\rho(\boldsymbol{A}_{m})$ is the spectral radius of $\boldsymbol{A}_{m}$, m=1,,Mm=1,...,M.
Condition B.1 states that the usual stability condition is satisfied by each of the regimes. It is necessary for Assumption B.1(I), as $\max(\rho(\boldsymbol{A}_{1}),...,\rho(\boldsymbol{A}_{M}))\leq\rho(\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\})$ (see Kheifets and Saikkonen, 2020, and the references therein). After estimation, Assumption B.1(I) can be checked for the solutions of interest.
Several methods have been proposed for computing bounds for the JSR, many of which are discussed in Chang and Blondel (2013). The accompanying R package sstvars (Virolainen, 2025a) implements the branch-and-bound method of Gripenberg (1996), but the implementation is computationally very demanding if the matrices are large and a relatively tight bound is required. The JSR toolbox in MATLAB (Jungers, 2023), in turn, automatically combines various methods to substantially enhance computational efficiency.
Appendix C Proofs
C.1 Proof of Lemma 1